# Collective Creation of Intimacy: Exploring the Cosplay Commission Practice within the Otome Game Community in China

### v2

@@ 4.3.2. Blurred Boundaries between Character Embodiment and Selfhood
Currently, there is no structured guideline for cos-commission. Cosplayers typically self-learn the personality of characters by watching anime, manga, and videos to provide an immersive and high-quality service. However, unlike other forms of paid companionship—such as sex workers, game companions (Shen et al., 2021), or host/hostess clubs in Japan (Takeyama, 2020)—cos-commission participants do not share a clear, unified understanding of their short-term relationship, which may lead to unexpected and unavoidable emotional transference. As such, it often becomes difficult for participants to distinguish whether their feelings are toward the character or the cosplayer themselves, leaving a paradoxical situation where the performance ends but the emotional immersion does not.
4.4. Platform-Related Vulnerabilities
@@
Digital platforms, as illustrated in Figure 1, are the main venues for organizing cos-commissions. However, they also present some vulnerabilities in the context of cos-commission.
4.4.1. Unregulated Information Exchange and Unwanted Contact
@@
On different platforms, participants have a variety of options to customize their media preferences. Video-sharing websites, such as BiliBili  [^5] BiliBili: A Chinese online video-sharing platform. https://bilibili.com/ [Last Accessed: May 17, 2025], provide open venues for users to demonstrate and disseminate positive experiences, and Xiaohongshu  [^6] Xiaohongshu: Known in English as RedNote, a Chinese social networking and e-commerce platform. https://xiaohongshu.com/ [Last Accessed: May 17, 2025] (RedNote), with a fusion of graphic and textual information, enables customers to search for cosplayers, comment on reviews and service authenticity. Douyin  [^7] Douyin: A short-video platform in mainland China and Hong Kong. While TikTok and Douyin share a similar interface, they operate separately. https://douyin.com/ [Last Accessed: May 17, 2025], which is called TikTok outside mainland China and Hong Kong, specializes in short videos, allowing customers to verify the skills of cosplayers via their previous performances. A second-hand trading platform, Xianyu  [^8] Xianyu: An online second-hand trading platform. https://goofish.com/ [Last Accessed: May 17, 2025], simplifies the cos-commission process with quick communication and payments and has many who are convenience-minded users. By using these various platforms, customers can find their preferred cosplayers and improve their cos-commission experiences with easy, safe, and high-quality transactions.
4.4.2. Privacy and Safety Concerns
@@
Although the benefits, there are also safety and privacy concerns over digital platforms. Professional cosplayers have raised apprehensions about being harassed or subjected to entanglements from a customer with whom they have built a relationship after a commission ends. For example, P10 reported that someone he knows achieved this: ”My friend posted a photo with her location on her birthday, and a customer followed the location to find her. She didn’t think the customer would, so she didn’t worry. … The customer arrived at her birthday party uninvited and insisted she drive him/her home, causing trouble for her and her girlfriend.” This is a case in point for a lack of anti-harassment and fair-use policies for paid companionship in real-world settings. Due to the feature of in-depth, dyadic communication and information exchange of cos-commission, cosplayers, and customers are pushed to a higher risk in Personally Identifiable Information (PII) disclosure.
@@
Female cosplayers also feared uncomfortable or mistaken encounters with male customers. Some commented that men are in general more physically powerful than women, and therefore if an accident were to happen they would have a harder time defending themselves. In cos-commission, if intimate interactions are to take place, participants often move to private spaces. Since there is no aid to be had from others, their anxiety level will increase even more.
@@ 5. Conclusion and Future work
In this preliminary study, we explored the motivations, experiences, and issues of the cos-commission participants within the Otome game community in China. Our results shows that cos-commission is a hybrid space in which the participants co-create meaningful connections. It is also a medium that allows its participants to rediscover themselves as well as find emotional recovery. However, the loose line between the performing roles and real intimacy may result in some uncomfortable effects, such as harmful behaviors driven by attachment and ”withdrawal symptoms” after performing.
Our research also reveals safety and privacy issues associated with the existing cos-commission ecosystem on digital platforms. Vulnerabilities, such as unwanted post-commission entanglements when PII (e.g., a phone number, or geolocation data) gets exposed or gender inequality among female cosplayers reflect system-level breakdowns in facilitating cross-platform intimate hybrid interactions. Strong safeguards, such as chat encryption, multi-factor identity verification, and adjustable privacy filters, are required to be put in place by the platforms. These could help to create safer spaces while maintaining the independent creativity of cos-commission culture.
Through our study of subculture communities as a case-in-point, we hope to contribute to the design of virtual world interaction at large. It would be interesting for future research to consider the longitudinal dynamics of cos-commission participation, including the transformation of emotional investments across multiple visits. Together with Otome game communities, we hope to co-design socio-technical infrastructures that align immersive intimacy with appropriate safeguards. If we (re)valorized emotional satisfaction alongside participant welfare, we might learn from subcultures how to construct a better digital world.

### v3

@@ 4.3.2. Blurred Boundaries
Currently, there is no structured guideline for cos-commission. Cosplayers typically self-learn the personality of characters by watching anime, manga, and videos to provide an immersive and high-quality service. However, unlike other forms of paid companionship, such as sex workers (Saha et al., 2024), paid game teammates (Shen et al., 2021), virtual lovers (Li et al., 2023) or host/hostess clubs in Japan (Takeyama, 2020), cos-commission participants do not share a clear, unified understanding of their short-term relationship, which may lead to unexpected and unavoidable emotional transference. As such, it often becomes difficult for participants to distinguish whether their feelings are towards the virtual character, the cosplayer, or their imaginations, leaving a paradoxical situation where the performance ends but the emotional immersion does not.
4.3.3. Unwanted Contact and Information Leakage
@@
Digital platforms, as illustrated in Figure 1, are the main venues for organizing cos-commissions. However, they also present some vulnerabilities.
@@
On different platforms, participants have a variety of options to customize their media preferences. Video-sharing websites, such as BiliBili  [^5] BiliBili: A Chinese online video-sharing platform. https://bilibili.com/ [Last Accessed: May 17, 2025], provide open spaces for users to showcase positive experiences. Xiaohongshu  [^6] Xiaohongshu: Known in English as RedNote, a Chinese social networking and e-commerce platform. https://xiaohongshu.com/ [Last Accessed: May 17, 2025] (RedNote), with a fusion of graphic and textual information, enables customers to search for cosplayers, comment on reviews and service authenticity. Douyin  [^7] Douyin: A short-video platform in mainland China and Hong Kong. While TikTok and Douyin share a similar interface, they operate separately. https://douyin.com/ [Last Accessed: May 17, 2025], which is called TikTok outside mainland China and Hong Kong, specializes in short videos, allows customers to verify the skills of cosplayers via their previous performances. Xianyu  [^8] Xianyu: An online second-hand trading platform. https://goofish.com/ [Last Accessed: May 17, 2025], a second-hand trading platform, simplifies the cos-commission process with quick communication and payments and has many who are convenience-minded users. By jumping back-and-forth between different platforms, customers can find their preferred cosplayers and improve their cos-commission experiences with easy, safe, and high-quality transactions.
@@
Despite the benefits, there are safety and privacy concerns over digital platforms. Professional cosplayers express worries about being harassed or subjected to entanglements from a customer with whom they have built a relationship after a commission ends. For example, P10 reported that someone he knows achieved this: “My friend posted a photo with her location on her birthday, and a customer followed the location to find her. She didn’t think the customer would, so she didn’t worry. … The customer arrived at her birthday party uninvited and insisted she drive him/her home, causing trouble for her and her girlfriend.” This case reflects a lack of anti-harassment mechanism and profile-management policy in cross-platform settings. Due to the feature of in-depth, dyadic communication and information exchange of cos-commission, cosplayers, and customers are pushed to a higher risk in personally identifiable information disclosure.
4.3.4. Gender-Related Vulnerability
@@
Speaking of gender, female cosplayers feared uncomfortable or mistaken encounters with male customers. Some commented that males are in general more physically powerful than females, and therefore if an accident were to happen they would have a harder time defending themselves. In cos-commission, if intimate interactions are to take place, participants sometimes move to private spaces. Since there is no aid to be had from others, their anxiety and vulnerability level will increase even more.
@@ 5. Conclusion and Future work
In this preliminary study, we explored the motivations, experiences, and issues of the cos-commission participants within the Otome game community in China. Our results shows that cos-commission provides a hybrid space in which customers co-create meaningful connections with cosplayers. It also serves as a medium that allows its participants to rediscover themselves as well as recover from bad emotional states. However, the ambiguous line between the performing roles and real intimacy may result in negative consequences, such as depression caused by attachment and “withdrawal symptoms” after performing.
Our investigation also reveals safety and privacy issues associated with the existing cos-commission practice. Vulnerabilities, such as unwanted post-commission entanglements when personally identifiable information (e.g., a phone number, or geolocation data) gets exposed or gender inequality among female cosplayers reflect system-level failures in facilitating cross-platform hybrid intimate interactions. Strong safeguards, such as chat encryption, multi-factor identity verification, and adjustable privacy filters, are required to be put in place by the platforms to help create safer spaces while maintaining the independent creativity of cos-commission culture.
Through our study of subculture communities as a case-in-point, we hope to contribute to the design of virtual world interaction at large. It would be valuable for future research to consider the longitudinal dynamics of cos-commission participation, including the transformation of emotional investments across multiple visits. Together with Otome game communities, we hope to co-design sociotechnical infrastructures that align immersive intimacy with appropriate safeguards. If we (re)valorized emotional satisfaction alongside participant welfare, we might learn from subcultures how to construct a better digital world.